OpenStudy (anonymous):

can some one help me please If the patient has to be administered a dosage of 2 tablets every 6 hours for 7 days, what is the number of tablets required for the prescribed dosage? so i knw the the conversion factor that describes the number of tablets taken for every 6 h. is 2 tablets/6h and The dose has to be administered for a "full treatment" spanning 7 days. The conversion factor that determine the number days in a full treatment is 7 days/1full treatment but now idk what to do?

OpenStudy (amistre64):

how many hours are in 7 days?

OpenStudy (amistre64):

7*24 right?

OpenStudy (amistre64):

how many 6 hour periods fit into that? 7*24/6 and with 2 tablets every 6 hours ... 2*7*24/6 tablets

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

is it 56?

OpenStudy (amistre64):

i believe so

OpenStudy (amistre64):

24/6 = 4 4*2 = 8, 8*7 = 56 yep

OpenStudy (anonymous):

thank you again